OpenStudy (anonymous):

URGENT PLEASE !! What is the mass, in grams, of a sample of 6.98 × 10^24 atoms of magnesium (Mg)? Show your work or explain the steps that you used to determine your answer.

OpenStudy (anonymous):

@dan815 @Hero @larryboxaplenty @pooja195

OpenStudy (dan815):

you know the mass it says for Mg on the periodic table?

OpenStudy (anonymous):

No.. I dont really know anything about this. Im not good at chemistry. I just started the course.

OpenStudy (anonymous):

find out how many atoms are in 1 mole of Mg, than convert that into grams

OpenStudy (anonymous):

Ex - Hydrogen has an atomic mass of 1.008, and therefore 1.008 moles of Hydrogen has 6.022 *10^22 atoms

OpenStudy (anonymous):

** If I remember correctly

OpenStudy (anonymous):

Atomic mass if slightly different dependent on your periodic table (google says something slightly different than your table usually does - usually with more sig figs)

OpenStudy (anonymous):

Lets say you have 3.45*10^24 atoms of hydrogen - how many grams is that?

OpenStudy (anonymous):

does 6.98 x 10^24 = 28.2 g

OpenStudy (anonymous):

\[3.45*10^{24} atoms \times \frac{ 1 mole H }{ 6.022*10^{24} atoms } \times \frac{ 1.008g }{ 1mole H }\]

OpenStudy (anonymous):

was does it say the weight it for Mg?

OpenStudy (anonymous):

*it should be 6.022*10^23

OpenStudy (anonymous):

I need to leave - I need to continue studying botany. If my math was wrong - I'm sure someone will say something.

OpenStudy (anonymous):

*math was for my example only